Carolyn is 25 and Vicky is 4. In how many years will Carolyn be four times as old as Vicky?
Answer by ankor@dixie-net.com(22740)

Carolyn is 25 and Vicky is 4.
In how many years will Carolyn be four times as old as Vicky?
:
Let y = no. of years for this to be true
:
25 + y = 4(4 + y)
25 + y = 16 + 4y
25 - 16 = 4y - y
9 = 3y
y = 3 years. C's age will be 4 times V's age
;
;
Check it:
25 + 3 = 4(4+3)